Tohban Report 2016-12-28

From RHESSI Wiki

Jump to: navigation, search


Tohban Reports
Start Date: 21 December 2016
End Date: 28 December 2016
Tohban: Martin Fivian
Tohban email: mfivian@ssl.berkeley.edu
Next Tohban: Jim McTiernan
List all reports



Contents

Solar Activity

The Sun was very quiet with a goes base level around B8. Some B-flare occurred on Dec 26 and Dec 27.


Memory Management

Decimation was quiet/vigorous for the full time period. The detectors are producing more counts without decimating than we can downlink, given all BGS passes are at night. In order not to decimate during Alma co-observations, we moved the read pointer from 2016-354-15:34 to 355-12:30 UTC. Alma co-observations start at 355-12:48 (Dec 20).

In an effort to reduce background counts, we lowered HV on det. 9 which didn't appear to have a significant effect. We further raised the slow valid threshold on 3F and 9F.

Detectors 3 and 9 are recording events during spacecraft day, and for 4 minutes prior to and after daylight.

Spacecraft Status

Detectors 3 and 9 are recording events. The cold tips are Cryocooler power is steady at 80W. Cryocooler transients are still frequent.


Data Gaps

skipped over data recorded between 2016-354-15:34 and 355-12:30 UTC.

Detector issues

Detectors 3 and 9 are powered on, and are recording events. Background rates are on the high side, between 1200 and 2000. In order to bring down the background rates, the following actions have been taken:

Dec 21: HV Det.9 lowered from 3000 to 2500V. However, the slow valid rates didn't significantly improve.

 2016-356-21:20:31 /ihvdac detector=9, voltage=129 ; 2500 V

Dec 22: The threshold for G3 Front slow was increased from 0x0C to 0x15, which brought the slow valid counts down.

 2016-357-20:52:39 /IDPULOAD VALUE=0x15 ; was 0x0C

Dec 26: The G9 front slow threshold was increased from 0x15 to 0x1C.

 2016-361-17:19:30 /IDPULOAD VALUE=0x1C ; was 0x15


Notes

Dec 21: The attenuator thresholds have been adjusted according to a reference background dead time of 14.5%. The thresholds are now as follows.

     Max A0  19.61
     Min A1  16.08
     Max A1  21.57
     Min A3  16.86

Spacecraft Management

Decimation Very Quiet/Vigorous
HLAT Decimation Rear decimation
Night time data (fronts) 4min of data
Night time data (rears) 4min of data
Require extra passes? No
Requirement for moving pointer? Yes
Attenuator operation Working as expected
Detector problems? See notes above.
Personal tools